### v1

# Dongseok Kim    Hyoungsun Choi    Mohamed Jismy Aashik Rasool    Gisung Oh

Abstract
We propose $\phi$-test, a global feature-selection and significance procedure for black-box predictors that combines Shapley attributions with selective inference. Given a trained model and an evaluation dataset, $\phi$-test performs SHAP-guided screening and fits a linear surrogate on the screened features via a selection rule with a tractable selective-inference form. For each retained feature, it outputs a Shapley-based global score, a surrogate coefficient, and post-selection $p$-values and confidence intervals in a global feature-importance table. Experiments on real tabular regression tasks with tree-based and neural backbones suggest that $\phi$-test can retain much of the predictive ability of the original model while using only a few features and producing feature sets that remain fairly stable across resamples and backbone classes. In these settings, $\phi$-test acts as a practical global explanation layer linking Shapley-based importance summaries with classical statistical inference.
Machine Learning, ICML
1 Introduction
Modern machine learning practice is dominated by high-performing black-box predictors on heterogeneous tabular features. In many scientific and engineering domains, these models are not merely auxiliary tools but primary objects of analysis: practitioners want to know not only how well a model predicts, but also which input features matter globally and how strongly they affect the output. Shapley-value-based attribution methods—most prominently SHAP—have become a de facto standard for this purpose, offering local explanations and global summaries via aggregates of absolute SHAP values. In practice, these summaries are mostly used as ranking heuristics: one applies a threshold or top-$K$ cutoff and reports the resulting list as ”important”, typically without formal uncertainty quantification, without correcting for the data-driven selection step, and without checking whether the selected features admit a coherent statistical effect description.
Classical statistics offers a complementary viewpoint. In linear and generalized linear models, feature effects are summarized in compact statistical tables that report coefficients, $p$-values, and confidence intervals, enabling hypothesis testing and effect-size interpretation at a glance. Domain experts continue to rely on such tables even when prediction uses a more flexible backbone. Bridging these worlds is nontrivial: regressing the outcome $y$ on raw features $x$ ignores the fitted black-box $f$, while regressing $f(x)$ on $x$ and performing variable selection induces post-selection bias. Recent Shapley-based variable-importance methods and SHAP-based hypothesis tests address global importance and, to some extent, uncertainty, but they typically do not model the selection step explicitly. As a result, they do not yield a post-selection global feature-importance table with coefficients, $p$-values, and confidence intervals for the retained features.
We propose the $\phi$-test (phi-test), a global explanation framework that turns SHAP values into statistically disciplined feature-importance tables for arbitrary black-box models. We first perform SHAP-based global screening to obtain a manageable candidate set of features, then fit a linear surrogate that regresses the black-box output (or class logit) on these candidates and applies a tractable selection procedure, and finally use selective inference to compute post-selection $p$-values and confidence intervals for the selected coefficients. The resulting global feature-importance table reports, for each feature, a SHAP-based global score, a surrogate coefficient estimate, and selection-adjusted uncertainty summaries, yielding an ablation-free global explanation with explicit statistical uncertainty.
Our contributions are as follows:
  • We introduce the $\phi$-test, combining SHAP-based global screening, surrogate regression, and selective inference to provide feature-wise post-selection significance for arbitrary black-box models, summarized in a global feature-importance table.
  • We specify the selection event, the relevant linear contrasts, and the resulting truncated-normal maps that power our global feature-importance table, building on the general theory of polyhedral selective inference.
  • We propose a real-data evaluation protocol based on performance retention, sparsity, and selection stability, and show that the $\phi$-test attains competitive predictive performance while selecting smaller and more stable or at least comparable feature sets than Shapley-based baselines. In addition, we summarize regression-based statistical support via a global feature-importance table with selective $p$-values and confidence intervals.
2 Related Work
From Local Attributions to Global Additive Explanations.
Several works lift instance-wise attributions to global additive descriptions of a predictor by decomposing $f$ into low-dimensional main and interaction components and showing that tools such as partial dependence and interventional Shapley values arise as linear functionals of this decomposition (hiabu2023unifying; fumagalli2024unifying). Other lines aggregate variable-importance measures or local explanations across Rashomon sets, datasets, or examples, treating them as distributions or geometric objects in an ”explanation space” and using them to explain distribution shifts or recourse patterns rather than a single model-specific ranking (donnelly2023rashomon; babbar2024different; xenopoulos2022gale; alfeo2023local). These methods demonstrate that local explanations can be turned into global effect summaries, but they typically provide point estimates of importance and do not attach selection-adjusted uncertainty to feature-wise statistics.
Global Variable-Importance Measures with Uncertainty Quantification.
A line of work formalizes global variable importance as a population-level contrast in oracle predictiveness between models that use all features and models that exclude a given subset, and derives efficient estimators whose asymptotic normality enables confidence intervals and hypothesis tests even with flexible learners (williamson2023general; williamson2020efficient). Building on this perspective, decorrelated importance parameters mitigate the impact of feature correlation while still supporting valid uncertainty quantification in high-dimensional or genomics applications via Shapley-style loss decompositions and bootstrap procedures (verdinelli2024decorrelated; johnsen2023inferring). Additional methods develop minipatch ensembles, simultaneous confidence sets for feature rankings, and conditional permutation schemes with formal type I error control for variable inclusion decisions (gan2022model; neuhof2024confident; chamma2023statistically). These approaches provide principled uncertainty quantification for pre-specified global importance functionals, but they do not explicitly model a feature-selection step or adjust their intervals and $p$-values for data-driven selection events.
Attribution-Driven Hypothesis Testing and Feature Screening.
Shapley-based attributions have been linked to conditional independence testing, for example in SHAP-XRT, which embeds local Shapley scores into a conditional randomization framework to obtain finite-sample valid tests for feature relevance (teneggi2022shap). Other approaches design attribution-driven screening rules in which Shapley scores are compared against data-driven thresholds or causal criteria, yielding adaptive selection procedures such as ShapHT plus and Causal SHAP that control false discoveries or preserve causal interpretability (yin2022adaptive; lamsaf2025causal). Complementary work studies the statistical significance of entire attribution rankings, proposing multiple-hypothesis tests that retrospectively certify the correctness of the top-$K$ features identified by an importance estimator (goldwasser2024statistical). These lines of work use attribution scores as inputs to hypothesis tests or screening rules, but they typically focus on local conditional independence decisions or on validating a fixed ranking, rather than producing a global feature-importance table with post-selection uncertainty for each selected feature.
Stability and Cross-Model Robustness in Feature Selection.
Several works modify local post hoc explainers to improve the stability of attributions under small input perturbations, for example by redesigning the sampling and weighting scheme of LIME in GLIME or by leveraging invariant risk minimization ideas to obtain explanations that vary smoothly across nearby examples (tan2023glime; dhurandhar2023locally). Other methods explicitly target uncertainty and stability in explanations, introducing Bayesian variants of attribution methods with credible intervals, learning global surrogate explanations that remain faithful under shifts, and designing rule-based or counterfactual objectives and evaluation metrics for stability in tabular decision problems (slack2021reliable; lakkaraju2020robust; guidotti2024stable; sepulveda2025enhancing). Cross-model robustness is studied by aggregating explanations across committees of models and using cross-model consensus scores to capture features that are relied on consistently across architectures (li2023cross). These efforts highlight the importance of stability and cross-model agreement, which we also target in our empirical evaluation, but they do not provide selective-inference-style uncertainty for feature-wise effects.
Post-Selection Inference for Regression-Based Feature Selection.
The post-selection inference literature develops tools for valid regression inference after data-driven feature selection by conditioning on the selection event, often exploiting the fact that procedures such as the lasso, forward stepwise, and least-angle regression induce polyhedral constraints on the response so that linear contrasts of the coefficients follow truncated normal laws (lee2016exact; tibshirani2016exact). Within a more general exponential-family framework, selective tests and confidence intervals are constructed to be most powerful and unbiased under arbitrary model-selection rules, providing a principled alternative to naive plug-in $t$-tests that ignore selection (fithian2014optimal; taylor2015statistical). Subsequent extensions introduce randomized selection steps, approximate Bayesian selective-likelihood formulations, and knockoff filters that address high-dimensional variable selection via synthetic controls and finite-sample false-discovery guarantees (tian2018selective; panigrahi2021integrative; barber2019knockoff). Together, these approaches show how to attach valid $p$-values and confidence intervals to coefficients in regression models chosen by adaptive feature selection, but they are typically developed for generic supervised learning tasks and do not directly address Shapley-guided global feature-importance tables for black-box predictors.
3 Method
We assume a trained black-box predictor $f:\mathcal{X}\to\mathbb{R}$ (or a real-valued score such as a class logit) and an evaluation dataset
$D=\{(x_{i},y_{i})\}_{i=1}^{n},\qquad x_{i}\in\mathbb{R}^{p}.$
Given $(f,D)$, the $\phi$-test produces a global feature-importance table whose rows index features $j=1,\dots,p$ and whose columns report: (i) a SHAP-based global score $I_{j}$; (ii) a surrogate regression coefficient $\widehat{\beta}_{j}$; and (iii) post-selection $p$-values and confidence intervals for features retained after a SHAP-guided selection step. Algorithm 1 summarizes the procedure.
We write $\mathcal{T}_{M}((I_{j})_{j=1}^{p})$ for the indices of the $M$ largest entries of $(I_{j})_{j=1}^{p}$, and
$(S,\widehat{\beta}_{S},\widehat{\sigma}^{2},A,b,\{v_{j}\}_{j\in S})=\mathcal{S}(X_{S_{0}},y^{(\mathrm{sur})})$
for the outputs of the chosen selection procedure (e.g., lasso or forward stepwise) applied to the screened design $X_{S_{0}}$ and surrogate response $y^{(\mathrm{sur})}$, where $(A,b)$ encode the corresponding polyhedral selection event and $v_{j}$ are contrast vectors for the selected coefficients.
3.1 Stage 1: SHAP-based global screening
Given a SHAP engine appropriate for $f$ (e.g., TreeSHAP or KernelSHAP), we compute local Shapley attributions
$\Phi\;=\;(\phi_{j}(x_{i}))_{i=1,\dots,n;\,j=1,\dots,p},\qquad\Phi_{ij}=\phi_{j}(x_{i}),$
and aggregate them into global scores
$I_{j}\;=\;\frac{1}{n}\sum_{i=1}^{n}\bigl|\Phi_{ij}\bigr|,\qquad j=1,\dots,p,$
though other norms (e.g., root-mean-square) could be used. We retain the top $M$ features according to $(I_{j})$,
$S_{0}\;=\;\mathcal{T}_{M}\bigl((I_{j})_{j=1}^{p}\bigr),$
where $M$ is a user- or budget-controlled parameter. Features outside $S_{0}$ keep their $I_{j}$ in the final table but are excluded from surrogate modelling and selective inference. In practice $M$ can be chosen moderately larger than a target sparsity level (e.g., between $K$ and $3K$), so that clearly unimportant features are discarded while leaving room for the selection step to refine $S_{0}$.
3.2 Stage 2: Surrogate regression and selection
We approximate how $f$ responds to changes in the screened features via a low-dimensional linear surrogate. For regression tasks we set
$y_{i}^{(\mathrm{sur})}=f(x_{i}),\qquad i=1,\dots,n,$
and for classification we take a scalar score for a chosen class (e.g., its logit or probability). Let $X\in\mathbb{R}^{n\times p}$ be the full design matrix and let $X_{S_{0}}$ denote its submatrix with columns indexed by $S_{0}$. We posit the working linear model
$y_{i}^{(\mathrm{sur})}=\beta_{0}+\sum_{j\in S_{0}}\beta_{j}x_{ij}+\varepsilon_{i},\qquad i=1,\dots,n.$
We view this as a linear approximation to the black-box predictions along the observed feature directions. The coefficients $\beta_{j}$ summarize how $f(x)$ varies when feature $j$ is perturbed within the support of $X_{S_{0}}$; they are not intended as causal effects, and their interpretability depends on the surrogate achieving a reasonable fit to $f$.
On $(X_{S_{0}},y^{(\mathrm{sur})})$ we apply a selection procedure whose selection event admits a polyhedral representation in $y^{(\mathrm{sur})}$, such as lasso with a fixed tuning parameter or forward stepwise selection with a fixed entry rule. We collect its outputs as
$(S,\widehat{\beta}_{S},\widehat{\sigma}^{2},A,b,\{v_{j}\}_{j\in S})=\mathcal{S}(X_{S_{0}},y^{(\mathrm{sur})}),$
where $S\subseteq S_{0}$ is the selected set, $\widehat{\beta}_{S}$ are refitted least-squares coefficients on $X_{S}$, $\widehat{\sigma}^{2}$ is a residual variance estimate, and
$\mathcal{E}=\{y^{(\mathrm{sur})}\in\mathbb{R}^{n}:Ay^{(\mathrm{sur})}\leq b\}$
encodes the selection event. For each $j\in S$, the contrast vector $v_{j}\in\mathbb{R}^{n}$ is chosen so that $v_{j}^{\top}y^{(\mathrm{sur})}$ is proportional to the least-squares estimate of $\beta_{j}$ under the model restricted to $S$ (e.g., via the usual OLS contrast construction).
3.3 Stage 3: Selective inference for surrogate coefficients
To correct for the data-driven selection in Stage 2 (Subsection 3.2), we adopt the Gaussian working model
$y^{(\mathrm{sur})}\sim\mathcal{N}\!\bigl(X_{S_{0}}\beta,\,\sigma^{2}I_{n}\bigr)$
and apply the truncated-normal selective-inference machinery from Section 4. For each $j\in S$ we define the linear contrast
$T_{j}=v_{j}^{\top}y^{(\mathrm{sur})},$
whose conditional law under the polyhedral selection event $\mathcal{E}$ is univariate Gaussian truncated to an interval determined by $(A,b)$. Using this truncated normal law with $\sigma^{2}$ replaced by $\widehat{\sigma}^{2}$, we compute a selective $p$-value and confidence interval for $\beta_{j}$ and write
$\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)=\bigl(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j}\bigr)$
for this one-dimensional selective-inference map.
3.4 Global feature-importance table
The $\phi$-test outputs a table indexed by features $j=1,\dots,p$. For each feature we report: (i) the SHAP-based global score $I_{j}$; (ii) the surrogate coefficient estimate $\widehat{\beta}_{j}$ if $j\in S$ (blank otherwise); and (iii) the selective $p$-value $p_{j}^{\mathrm{sel}}$ and confidence interval $\mathrm{CI}_{j}$ if $j\in S$. Features with large $I_{j}$ but $j\notin S$ are globally influential under SHAP yet not supported by the surrogate at the chosen selection threshold, whereas features in $S$ with modest $I_{j}$ correspond to more subtle but statistically stable effects. By presenting these quantities in a single table, the $\phi$-test provides an ablation-free, global explanation of how the black-box depends on its inputs and attaches post-selection uncertainty to each selected feature.
Algorithm 1 $\phi$-test
0:  black-box $f$, data $(x_{i})_{i=1}^{n}$, SHAP engine, budget $M$
0:  $\{I_{j},\widehat{\beta}_{j},p_{j}^{\mathrm{sel}},\mathrm{CI}_{j}\}_{j=1}^{p}$
1:  $\Phi\leftarrow(\phi_{j}(x_{i}))_{i=1,\dots,n;\,j=1,\dots,p}$
2:  $I_{j}\leftarrow\frac{1}{n}\sum_{i=1}^{n}|\Phi_{ij}|\quad(j=1,\dots,p)$
3:  $S_{0}\leftarrow\mathcal{T}_{M}\bigl((I_{j})_{j=1}^{p}\bigr)$
4:  $y_{i}^{(\mathrm{sur})}\leftarrow f(x_{i})\quad(i=1,\dots,n)$
5:  $X_{S_{0}}\leftarrow$ submatrix of $X$ with columns in $S_{0}$
6:  $(S,\widehat{\beta}_{S},\widehat{\sigma}^{2},A,b,\{v_{j}\}_{j\in S})\leftarrow\mathcal{S}(X_{S_{0}},y^{(\mathrm{sur})})$
7:  for $j\in S$ do
8:  $T_{j}\leftarrow v_{j}^{\top}y^{(\mathrm{sur})}$
9:  $(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j})\leftarrow\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)$
10:  end for
11:  return table with $I_{j}$ for all $j$ and $(\widehat{\beta}_{j},p_{j}^{\mathrm{sel}},\mathrm{CI}_{j})$ for $j\in S$
A summary of the notation used in Algorithm 1 is provided in Appendix A. In our experiments (Section 5), we use a split-sample implementation in which SHAP-based selection and surrogate refitting are run on disjoint halves of the training data.
4 Theory
This section formalizes the statistical guarantees underlying the $\phi$-test. We work under a Gaussian surrogate model for the black-box predictions and a polyhedral representation of the selection event, and use standard selective-inference tools to characterize the conditional law of linear contrasts. Throughout, the black-box predictor $f$ and covariates $X$ are treated as fixed, and randomness arises only from the surrogate response.
4.1 Selective-inference setup
Recall the surrogate target and design from Section 3. Let $y^{(\mathrm{sur})}\in\mathbb{R}^{n}$ and $X_{S_{0}}\in\mathbb{R}^{n\times|S_{0}|}$ denote the surrogate response and the screened design matrix. We adopt the working model
$y^{(\mathrm{sur})}\sim\mathcal{N}\!\bigl(X_{S_{0}}\beta,\,\sigma^{2}I_{n}\bigr),$ (1)
for some unknown $\beta\in\mathbb{R}^{|S_{0}|}$ and $\sigma^{2}>0$. The $\phi$-test applies a variable-selection procedure to $(X_{S_{0}},y^{(\mathrm{sur})})$ and collects its outputs as
$(S,\widehat{\beta}_{S},\widehat{\sigma}^{2},A,b,\{v_{j}\}_{j\in S})\;=\;\mathcal{S}\bigl(X_{S_{0}},y^{(\mathrm{sur})}\bigr),$
where $S\subseteq S_{0}$ is the selected set, $\widehat{\beta}_{S}$ are refitted least-squares coefficients on $X_{S}$, $\widehat{\sigma}^{2}$ is a residual variance estimate, and $(A,b)$ encode the selection event.
We assume that the selection method admits a polyhedral representation in $y^{(\mathrm{sur})}$: there exist $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$ such that
$\mathcal{E}=\mathcal{E}(S)=\{y^{(\mathrm{sur})}\in\mathbb{R}^{n}:Ay^{(\mathrm{sur})}\leq b\},$ (2)
which encodes the event that the procedure selects $S$ (possibly together with a sign pattern or entry order).
For a fixed selected set $S$, let
$\widehat{\beta}_{S}=(X_{S}^{\top}X_{S})^{-1}X_{S}^{\top}y^{(\mathrm{sur})}$
denote the ordinary least-squares estimator in the surrogate model restricted to $S$. For each $j\in S$, we are interested in testing
$H_{0,j}:\beta_{j}=0$
and constructing a confidence interval for $\beta_{j}$, conditional on the event $\mathcal{E}$ that $S$ has been selected.
4.2 Truncated-normal law and selective validity
Let $v_{j}\in\mathbb{R}^{n}$ be a contrast vector such that
$T_{j}\;=\;v_{j}^{\top}y^{(\mathrm{sur})}$
is proportional to $\widehat{\beta}_{j}$; for example, when $S$ is fixed one may take
$v_{j}=e_{j}^{\top}(X_{S}^{\top}X_{S})^{-1}X_{S}^{\top},$
where $e_{j}$ is the canonical basis vector for feature $j$ within $S$. Writing $\mu=X_{S_{0}}\beta$ in (1), define
$\theta_{j}=v_{j}^{\top}\mu,\qquad\tau_{j}^{2}=\sigma^{2}\|v_{j}\|_{2}^{2}.$
Under (1) and (2), decompose
$T_{j}=v_{j}^{\top}y^{(\mathrm{sur})},\qquad W_{j}=P_{v_{j}^{\perp}}y^{(\mathrm{sur})},$
where $P_{v_{j}^{\perp}}$ is the orthogonal projector onto the subspace orthogonal to $v_{j}$. Then $(T_{j},W_{j})$ is jointly Gaussian with $T_{j}\sim\mathcal{N}(\theta_{j},\tau_{j}^{2})$ and $T_{j}$ independent of $W_{j}$. Conditioning on $W_{j}$ and restricting to the affine line $\{W_{j}+tv_{j}:t\in\mathbb{R}\}$, the event $\mathcal{E}=\{Ay^{(\mathrm{sur})}\leq b\}$ reduces to linear inequalities in $t$, so that $T_{j}$ is constrained to an interval $a_{j}(W_{j}),b_{j}(W_{j})]$. Consequently,
$T_{j}\mid(W_{j},\mathcal{E})\sim\mathcal{N}(\theta_{j},\tau_{j}^{2})\text{ truncated to }[a_{j}(W_{j}),b_{j}(W_{j})].$
This truncated normal law is the basis for the selective $p$-values and confidence intervals reported by the $\phi$-test. Under $H_{0,j}:\theta_{j}=0$, the conditional distribution of $T_{j}$ given $(W_{j},\mathcal{E})$ is a centered truncated normal, from which one can compute one- or two-sided tail probabilities and invert these to obtain selective confidence intervals. For notational convenience, we write
$\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)=\bigl(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j}\bigr)$
for the map that takes $T_{j}$ together with $(A,b)$ and $\widehat{\sigma}^{2}$, and returns the selective $p$-value and confidence interval computed from this truncated normal law.
Lemma 4.1 (Ancillarity of SHAP screening).
Let $S_{0}=\mathcal{T}_{M}\bigl((I_{j})_{j=1}^{p}\bigr)$ denote the candidate feature set produced by Stage 1 (Subsection 3.1) of the $\phi$-test based on global SHAP scores. Under (1), the set $S_{0}$ is non-random with respect to $y^{(\mathrm{sur})}$, so all probabilities can be interpreted conditionally on $(f,X,S_{0})$.
Proposition 4.2 (Selective validity).
Work under (1) and (2), and condition on $(f,X,S_{0})$ as in Lemma 4.1. Fix $j\in S$ and consider $H_{0,j}:\beta_{j}=0$. Let $p_{j}^{\mathrm{sel}}$ and $\mathrm{CI}_{j}$ denote the selective $p$-value and $(1-\alpha)$ confidence interval returned by $\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)$. Then, for all $\alpha\in[0,1]$,
$\mathbb{P}\bigl(p_{j}^{\mathrm{sel}}\leq\alpha\,\big|\,f,X,S_{0},j\in S\bigr)$ $\leq\alpha,$
$\mathbb{P}\bigl(\beta_{j}\in\mathrm{CI}_{j}\,\big|\,f,X,S_{0},j\in S\bigr)$ $\geq 1-\alpha.$
That is, the selective $p$-values control conditional type-I error and the intervals achieve at least nominal conditional coverage for the surrogate coefficients given that feature $j$ has been selected.
Proof sketch.
The result follows from the truncated-normal characterization of $T_{j}\mid(W_{j},\mathcal{E})$ under polyhedral selective inference, by computing tail probabilities under $H_{0,j}$ and inverting the corresponding tests. ∎
4.3 Application to the $\phi$-test
The $\phi$-test instantiates this framework as follows. Stage 1 (Subsection 3.1) (SHAP-based screening) defines
$S_{0}=\mathcal{T}_{M}\bigl((I_{j})_{j=1}^{p}\bigr)$
from global SHAP scores; conditional on the realized $S_{0}$, the surrogate model (1) is defined on $X_{S_{0}}$. Stage 2 (Subsection 3.2) applies a selection procedure on $(X_{S_{0}},y^{(\mathrm{sur})})$ whose selection event is polyhedral, yielding
$(S,\widehat{\beta}_{S},\widehat{\sigma}^{2},A,b,\{v_{j}\}_{j\in S})=\mathcal{S}(X_{S_{0}},y^{(\mathrm{sur})}).$
For each $j\in S$, Stage 3 (Subsection 3.3) constructs a contrast $v_{j}$ corresponding to $\widehat{\beta}_{j}$ in the refitted least-squares model on $X_{S}$, defines $T_{j}=v_{j}^{\top}y^{(\mathrm{sur})}$, and evaluates the truncated normal law of $T_{j}\mid(W_{j},\mathcal{E})$ with $\sigma^{2}$ replaced by $\widehat{\sigma}^{2}$. The resulting selective $p$-values and confidence intervals are summarized as
$(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j})=\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr),$
and Proposition 4.2 describes their conditional validity.
Remark 4.3 (Scope and interpretation).
The guarantees above are conditional on the screening set $S_{0}$, the design $X_{S_{0}}$, and the fitted black-box $f$, and apply to the surrogate coefficients that summarize how $f$ responds along the observed feature directions. We treat $S_{0}$ as fixed and do not claim causal or interventional validity; the goal is a post-selection-corrected global feature-importance summary of the fitted black-box.
When the selection and inference data are independent, as in the split-sample implementation in Section 5, $T_{j}$ obeys the usual Gaussian linear-model law on the inference sample. Classical $t$-tests are then valid, and truncated-normal corrections are only needed when selection and inference share observations.
Remark 4.4 (Split-sample $\phi$-test).
In our empirical evaluation (Section 5), we use a split-sample version of $\phi$-test: one half of the training data is used for SHAP-guided screening and surrogate selection, and the other half for refitting and inference. Because these splits are independent, classical $t$-tests on the inference split are already valid post selection and can be viewed as a conservative instance of the truncated-normal framework above.
5 Experiments
We evaluate $\phi$-test on real tabular regression problems with pre-trained black-box predictors. We ask how well the selected features retain predictive performance, how sparse and stable the explanations are, and how robust the selected feature sets are to changing the backbone model class.
In all experiments we use a split-sample $\phi$-test: one half of the training data for SHAP-based selection, the other for refitting and $t$-tests, which are valid by independence and form a conservative special case of our truncated-normal framework.
5.1 Experimental setup
For each dataset we train a black-box predictor $f$ on the full feature set using two backbones, XGBoost and a feed-forward multilayer perceptron (MLP). We split the data into training, validation, and test sets; models are trained with early stopping on the validation split, and backbone hyperparameters are tuned by a small grid search on the training/validation data only. Once trained, $f$ is held fixed and serves as the common object of explanation for all methods.
Given $f$ and the selection split $\{x_{i}\}$ of the training data, we compute per-sample Shapley attributions $\phi_{j}(x_{i})$ using TreeSHAP for XGBoost and KernelSHAP for MLP. All competing methods share these precomputed Shapley values on the same selection split. For $\phi$-test, we define the surrogate response $y^{(\mathrm{sur})}_{i}=f(x_{i})$ and apply SHAP-based global screening followed by linear surrogate regression and selective inference as in Section 3; the baselines consume the same attributions but aggregate or threshold them without a selective-inference step. To quantify selection stability, we repeat the full pipeline (training, explanation, evaluation) over multiple random resamples of the training data.
We evaluate on four real-valued regression benchmarks: Concrete (compressive strength prediction with $p=8$ physical inputs), AirQuality (gas sensor measurements predicting an air-quality index), EnergyEfficiency (building energy loads from eight geometric and material attributes), and kin8nm (kinematic response as a function of eight physical variables).
5.2 Compared methods
We compare $\phi$-test against four Shapley-based baselines for global feature selection:
  • SHAP-TopK. (lundberg2017unified) Rank features by $I_{j}=\frac{1}{n}\sum_{i=1}^{n}|\phi_{j}(x_{i})|$ and select the top $K$ features.
  • SPVIM-Boot. (williamson2020efficient) Shapley-based variable importance with bootstrap confidence intervals; select features whose intervals exclude zero.
  • SHAP-HT. (yin2022adaptive) Marginal SHAP-based hypothesis tests of feature relevance with multiple-testing correction.
  • StableSHAP. (goldwasser2024statistical) A stability-oriented aggregation that re-fits the backbone on bootstrap samples and selects features that appear frequently among the top-ranked SHAP features.
All methods are constrained to select at most $K$ features, with $K=5$ in our main experiments to emphasize sparsity; SPVIM-Boot may return a larger set when its confidence intervals deem many variables significant.
5.3 Evaluation criteria
We report four scalar criteria for performance, sparsity, stability, and cross-backbone robustness. Let $\hat{S}$ denote the selected feature set and let $\mathrm{Perf}(\cdot)$ be a test performance metric (we use $R^{2}$ in all experiments) of a linear model fitted using only features in $\hat{S}$.
Fidelity.
$\mathrm{Fidelity}=\frac{\mathrm{Perf}(\hat{S})}{\mathrm{Perf}(\text{full})}\times 100\%.$ (3)
Sparsity.
[…]
Stability.
From $R$ repeated random train–test splits, we rerun the entire selection pipeline and obtain selected feature sets $\hat{S}^{(1)},\dots,\hat{S}^{(R)}$ for each method; we then compute the average pairwise Jaccard similarity
$\mathrm{Stability}=\frac{2}{R(R-1)}\sum_{r<s}\frac{|\hat{S}^{(r)}\cap\hat{S}^{(s)}|}{|\hat{S}^{(r)}\cup\hat{S}^{(s)}|}.$ (5)
Robustness.
For XGBoost and MLP backbones, let $\hat{S}_{\mathrm{XGB}}$ and $\hat{S}_{\mathrm{MLP}}$ be the selected sets on the same dataset:
$\mathrm{Robustness}=\frac{|\hat{S}_{\mathrm{XGB}}\cap\hat{S}_{\mathrm{MLP}}|}{|\hat{S}_{\mathrm{XGB}}\cup\hat{S}_{\mathrm{MLP}}|}.$ (6)
5.4 Quantitative results
Tables 14 summarize Fidelity, Sparsity, Stability, and Robustness across the four regression benchmarks. On Concrete and AirQuality (Tables 1 and 2), all Shapley-based methods except SHAP-HT achieve high Fidelity; SPVIM-Boot is sometimes slightly higher but uses more features, whereas $\phi$-test attains comparable Fidelity with five features and perfect cross-backbone Robustness, while SHAP-TopK and StableSHAP show lower Robustness and, on AirQuality, reduced Stability. On EnergyEfficiency (Table 3), $\phi$-test, SHAP-TopK, SPVIM-Boot, and StableSHAP have similar high Fidelity, with SPVIM-Boot using 6–7 features and $\phi$-test maintaining competitive performance with five. On kin8nm (Table 4), the non-degenerate methods are perfectly stable and backbone-robust; SPVIM-Boot reaches the highest Fidelity with all eight features, while $\phi$-test consistently improves on SHAP-TopK at the same sparsity level.
Table 1: Summary of feature selection performance on the Concrete dataset.
Model Method Fidelity (%) Sparsity Stability Robustness
XGBoost $\phi$-test 65.20 5 1.00 1.00
SHAP-TopK 65.20 5 0.93 0.67
SPVIM-Boot 66.57 7 1.00 0.88
SHAP-HT 0.00 0 0.80 0.00
StableSHAP 64.78 5 0.96 0.67
MLP $\phi$-test 82.36 5 0.93 1.00
SHAP-TopK 85.10 5 1.00 0.67
SPVIM-Boot 84.51 8 0.88 0.88
SHAP-HT -0.81 1 0.26 0.00
StableSHAP 77.83 5 0.83 0.67
Table 2: Summary of feature selection performance on the AirQuality dataset.
Model Method Fidelity (%) Sparsity Stability Robustness
XGBoost $\phi$-test 99.88 5 0.93 1.00
SHAP-TopK 99.88 5 0.88 0.67
SPVIM-Boot 100.21 12 1.00 1.00
SHAP-HT 24.57 2 0.52 0.00
StableSHAP 99.88 5 0.78 0.43
MLP $\phi$-test 98.62 5 0.89 1.00
SHAP-TopK 98.34 5 0.54 0.67
SPVIM-Boot 98.86 12 1.00 1.00
SHAP-HT 0.00 0 0.47 0.00
StableSHAP 96.47 5 0.37 0.43
Table 3: Summary of feature selection performance on the EnergyEfficiency dataset.
Model Method Fidelity (%) Sparsity Stability Robustness
XGBoost $\phi$-test 88.99 5 0.87 0.43
SHAP-TopK 88.99 5 1.00 0.43
SPVIM-Boot 88.96 6 1.00 0.63
SHAP-HT 36.49 1 0.35 1.00
StableSHAP 90.78 5 1.00 0.43
MLP $\phi$-test 94.68 5 0.76 0.43
SHAP-TopK 94.58 5 0.93 0.43
SPVIM-Boot 95.03 7 0.84 0.63
SHAP-HT 41.70 1 0.17 1.00
StableSHAP 94.73 5 0.65 0.43
Table 4: Summary of feature selection performance on the kin8nm dataset.
Model Method Fidelity (%) Sparsity Stability Robustness
XGBoost $\phi$-test 49.19 5 1.00 1.00
SHAP-TopK 47.84 5 1.00 1.00
SPVIM-Boot 51.46 8 1.00 1.00
SHAP-HT 0.00 0 0.00 0.00
StableSHAP 7.70 5 1.00 1.00
MLP $\phi$-test 43.02 5 1.00 1.00
SHAP-TopK 41.73 5 1.00 1.00
SPVIM-Boot 45.04 8 1.00 1.00
SHAP-HT 41.35 4 0.00 1.00
StableSHAP 6.40 5 1.00 1.00
5.5 Example global selective inference table
Table 5 shows the global feature-importance table produced by $\phi$-test on AirQuality with an XGBoost backbone. For the strongest pollutants, SHAP scores, regression coefficients, and very small selective $p$-values all indicate a positive and statistically well-supported effect, with selective 95% confidence intervals lying strictly above zero. In contrast, NO2(GT) has a moderate SHAP score but a selective $p$-value of $0.12$ and a confidence interval that crosses zero, suggesting that its effect is not conclusively different from zero once conditioning on the selection event. This illustrates how the table combines Shapley-based global importance with post-selection $p$-values and intervals to distinguish strongly supported features from borderline ones.
Table 5: Split-sample selective regression table for AirQuality ($\phi$-test, XGBoost backbone). Reported $p$-values and confidence intervals come from $t$-tests on the held-out inference split.
Feature SHAP Coef SE $z$ p-value 95% CI
C6H6(GT) 0.4487 0.0628 0.0085 7.4229 0.0000 [0.0462, 0.0795]
NOx(GT) 0.1717 0.0060 0.0005 12.2899 0.0000 [0.0050, 0.0069]
PT08.S1(CO) 0.0934 0.0008 0.0002 4.3770 0.0000 [0.0004, 0.0012]
NMHC(GT) 0.0820 0.0012 0.0002 7.1986 0.0000 [0.0008, 0.0015]
NO2(GT) 0.0600 0.0014 0.0009 1.5684 0.1178 [-0.0004, 0.0032]
Residual (unselected) 0.4707 [–, –]
5.6 Ablation study
We conduct a small ablation on the Concrete dataset with an XGBoost backbone, keeping the backbone and SHAP engine fixed and varying only the selection step among SHAP-guided lasso (SHAP + Lasso), plain lasso without SHAP screening (Lasso-only), SHAP-guided forward stepwise selection (SHAP + Stepwise), and a more strongly regularized lasso variant (Lasso-strong). As summarized in Table 6, all four variants achieve similar Fidelity and high Stability/Robustness, while SHAP-guided lasso provides a convenient default that balances sparsity and fidelity, with stronger regularization trading a small drop in Fidelity for a sparser explanation.
Table 6: Ablation on the selection step for the Concrete dataset (XGBoost backbone).
Method Fidelity (%) Sparsity Stability Robustness
SHAP + Lasso 66.3 5 1.00 1.00
Lasso-only 65.9 5 0.95 0.88
SHAP + Stepwise 64.0 4 0.93 0.82
Lasso-strong 62.0 3 0.96 0.90
6 Discussion and Limitations
Discussion.
Across our experiments (Tables 14), $\phi$-test generally attains predictive performance comparable to the strongest Shapley-based baselines while using substantially fewer features than dense Shapley-VIM procedures and yielding more stable and backbone-consistent selections than simple SHAP thresholding or stability-based heuristics. Rather than stopping at a global ranking, $\phi$-test turns Shapley attributions into an ANOVA-style feature table with post-selection $p$-values and confidence intervals: the example in Table 5 illustrates how this separates strongly supported drivers from borderline variables and corrects the overly optimistic uncertainty estimates obtained from naive regression on the same selected set. The ablation in Table 6 further suggests that these gains are not tied to a particular selection routine, as several SHAP-guided and plain selection rules exhibit similar predictive and sparsity profiles once embedded in the same surrogate-and-selective-inference layer. Overall, in the small-to-moderate dimensional tabular regression settings we study, the results support $\phi$-test as a practical way to add statistically principled, globally interpretable feature-importance tables on top of existing Shapley pipelines, without noticeably sacrificing the predictive fidelity of strong black-box explanations.
Limitations.
$\phi$-test relies on precomputed Shapley attributions (e.g., TreeSHAP or KernelSHAP), so its computational cost and robustness inherit the strengths and weaknesses of the underlying SHAP engine, especially for high-dimensional neural predictors. The linear surrogate keeps the global table easy to read but can only approximate nonlinear and interaction effects, so the resulting post-selection $p$-values and intervals are most informative when the surrogate provides a reasonable fit to $f$ and should be interpreted with caution otherwise. Our theoretical guarantees also rest on a Gaussian working model for the surrogate response and do not account for uncertainty in the training of $f$, so formal coverage statements should be read as properties of the fitted black-box rather than of the full data-generating process.
Future work.
Future directions include extending $\phi$-test to grouped features and low-order interactions, exploring richer yet interpretable surrogate families (such as generalized additive or generalized linear models) within the same selective-inference framework, and evaluating the method on larger and more structured domains where model classes are frequently swapped during development. It would also be natural to relax the Gaussian working model, incorporate uncertainty about the training of $f$ into the selective-inference layer, and investigate how $\phi$-test behaves in settings with many more features than observations or with strong domain-specific structural constraints.
7 Conclusion
We propose $\phi$-test, a global feature-selection and significance procedure for black-box predictors that combines Shapley attributions with selective inference to account for data-driven feature selection. Rather than stopping at a variable ranking, $\phi$-test outputs a global feature-importance table that, for each feature, reports a Shapley-based global importance score, a surrogate effect estimate, and post-selection uncertainty in a familiar regression-style format. On real tabular regression benchmarks with tree-based and neural backbones, $\phi$-test often retains much of the predictive ability of the underlying model while using only a few features and producing feature sets that remain fairly stable across resamples and backbone classes. We view $\phi$-test as a pragmatic link between Shapley-based explanations and classical statistical inference, and as a table-based summary that helps domain experts interpret and scrutinize complex predictive models.
Impact Statement
This work seeks to improve the transparency of black-box predictors in tabular domains by turning Shapley-based explanations into global feature-importance tables with selection-adjusted $p$-values and confidence intervals. By combining SHAP-guided screening, surrogate regression, and selective inference, $\phi$-test offers practitioners a compact, statistically principled summary of which features their models rely on and how strongly, without requiring ablation studies or changes to the underlying predictor. We expect this to be useful in scientific and engineering settings where stakeholders are already accustomed to interpreting classical regression summaries and statistical test tables.
The framework remains observational and model-based: the global tables describe how a fitted predictor responds to observed feature variations and do not provide causal guarantees. Misinterpreting these summaries as causal, or over-trusting selective $p$-values outside their assumptions, could lead to misguided decisions in high-stakes applications. We mitigate these risks by clearly stating the non-causal scope of the method, and we plan to release an open-source implementation so that users can inspect, stress-test, and adapt $\phi$-test to their own domains.
Appendix A Notation for Algorithm 1
Table 7 summarizes the main symbols used in Algorithm 1. For convenience, we collect here the notation that appears in the SHAP-guided screening, surrogate fitting, and selective-inference steps of the $\phi$-test. All quantities in the table are defined in the main text; this section is intended only as a quick reference.
Table 7: Notation used in Algorithm 1.
Symbol Description
$\Phi_{ij}$ SHAP value $\phi_{j}(x_{i})$
$I_{j}$ Global SHAP score $\frac{1}{n}\sum_{i=1}^{n}|\Phi_{ij}|$
$S_{0}$ Index set of top-$M$ features by $(I_{j})$
$\mathcal{T}_{M}((I_{j})_{j=1}^{p})$ Indices of the $M$ largest $I_{j}$
$y_{i}^{(\mathrm{sur})}$ Surrogate response $f(x_{i})$
$X_{S_{0}}$ Design matrix $X$ restricted to columns in $S_{0}$
$\mathcal{S}(X_{S_{0}},y^{(\mathrm{sur})})$ Selection map returning $(S,\widehat{\beta}_{S},\widehat{\sigma}^{2},A,b,\{v_{j}\}_{j\in S})$
$T_{j}$ Contrast $v_{j}^{\top}y^{(\mathrm{sur})}$ for feature $j$
$\mathcal{I}_{\mathrm{TN}}(T_{j};A,b,\widehat{\sigma}^{2})$ Truncated-normal map returning $(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j})$
Appendix B Details for Polyhedral Selective Inference
B.1 Truncated-normal law for linear contrasts
In this subsection we derive the truncated-normal characterization of the linear contrasts used by the $\phi$-test under the Gaussian surrogate model and a polyhedral selection event.
Recall the working model from (1),
$y^{(\mathrm{sur})}\sim\mathcal{N}\!\bigl(\mu,\,\sigma^{2}I_{n}\bigr),\qquad\mu=X_{S_{0}}\beta,$
and the polyhedral representation of the selection event from (2),
$\mathcal{E}=\mathcal{E}(S)=\{y^{(\mathrm{sur})}\in\mathbb{R}^{n}:Ay^{(\mathrm{sur})}\leq b\},$
for some matrix $A\in\mathbb{R}^{m\times n}$ and vector $b\in\mathbb{R}^{m}$. Fix a selected feature $j\in S$ and let $v_{j}\in\mathbb{R}^{n}$ be a contrast vector such that
$T_{j}=v_{j}^{\top}y^{(\mathrm{sur})}$
is proportional to the least-squares coefficient $\widehat{\beta}_{j}$ in the surrogate regression on $X_{S}$. We write
$\theta_{j}=v_{j}^{\top}\mu,\qquad\tau_{j}^{2}=\sigma^{2}\|v_{j}\|_{2}^{2}.$
Orthogonal decomposition.
Define the vector
$c_{j}=\frac{v_{j}}{\|v_{j}\|_{2}^{2}}\in\mathbb{R}^{n},$
so that $v_{j}^{\top}c_{j}=1$. We decompose $y^{(\mathrm{sur})}$ into a component along $v_{j}$ and an orthogonal remainder by
$y^{(\mathrm{sur})}=c_{j}T_{j}+W_{j},\qquad T_{j}=v_{j}^{\top}y^{(\mathrm{sur})},\qquad W_{j}=\bigl(I_{n}-c_{j}v_{j}^{\top}\bigr)y^{(\mathrm{sur})}.$ (7)
By construction, $W_{j}$ lies in the subspace orthogonal to $v_{j}$:
$v_{j}^{\top}W_{j}=v_{j}^{\top}(I_{n}-c_{j}v_{j}^{\top})y^{(\mathrm{sur})}=v_{j}^{\top}y^{(\mathrm{sur})}-(v_{j}^{\top}c_{j})(v_{j}^{\top}y^{(\mathrm{sur})})=0,$
so (7) is the usual decomposition of $y^{(\mathrm{sur})}$ into its projection onto $\mathrm{span}\{v_{j}\}$ plus the orthogonal remainder.
Since $y^{(\mathrm{sur})}\sim\mathcal{N}(\mu,\sigma^{2}I_{n})$, any linear transformation of $y^{(\mathrm{sur})}$ is Gaussian. In particular, $(T_{j},W_{j})$ is jointly Gaussian. The mean and variance of $T_{j}$ are
$\mathbb{E}[T_{j}]=v_{j}^{\top}\mu=\theta_{j},\qquad\mathrm{Var}(T_{j})=v_{j}^{\top}(\sigma^{2}I_{n})v_{j}=\sigma^{2}\|v_{j}\|_{2}^{2}=\tau_{j}^{2}.$
The covariance between $T_{j}$ and $W_{j}$ is
$\mathrm{Cov}(T_{j},W_{j})=\mathrm{Cov}\bigl(v_{j}^{\top}y^{(\mathrm{sur})},(I_{n}-c_{j}v_{j}^{\top})y^{(\mathrm{sur})}\bigr)=v_{j}^{\top}(\sigma^{2}I_{n})(I_{n}-c_{j}v_{j}^{\top})=\sigma^{2}\bigl(v_{j}^{\top}-v_{j}^{\top}c_{j}v_{j}^{\top}\bigr)=0,$
because $v_{j}^{\top}c_{j}=1$. Hence $T_{j}$ and $W_{j}$ are independent, with
$T_{j}\sim\mathcal{N}(\theta_{j},\tau_{j}^{2}),\qquad W_{j}\sim\mathcal{N}\bigl((I_{n}-c_{j}v_{j}^{\top})\mu,\,\sigma^{2}(I_{n}-c_{j}v_{j}^{\top})\bigr).$
Restriction of the selection event to an affine line.
We now express the polyhedral selection event $\mathcal{E}$ in the coordinates $(T_{j},W_{j})$. Let $a_{\ell}^{\top}$ denote the $\ell$-th row of $A$ and $b_{\ell}$ the corresponding entry of $b$, so that
$Ay^{(\mathrm{sur})}\leq b\quad\Longleftrightarrow\quad a_{\ell}^{\top}y^{(\mathrm{sur})}\leq b_{\ell}\quad\text{for all }\ell=1,\dots,m.$
Substituting the decomposition (7) gives
$a_{\ell}^{\top}(c_{j}T_{j}+W_{j})\leq b_{\ell}\quad\Longleftrightarrow\quad(a_{\ell}^{\top}c_{j})T_{j}\leq b_{\ell}-a_{\ell}^{\top}W_{j}.$
For each constraint, there are three cases:
  • If $a_{\ell}^{\top}c_{j}>0$, the inequality is equivalent to an upper bound
    $T_{j}\leq\frac{b_{\ell}-a_{\ell}^{\top}W_{j}}{a_{\ell}^{\top}c_{j}}.$
  • If $a_{\ell}^{\top}c_{j}<0$, the inequality is equivalent to a lower bound
    $T_{j}\geq\frac{b_{\ell}-a_{\ell}^{\top}W_{j}}{a_{\ell}^{\top}c_{j}}.$
  • If $a_{\ell}^{\top}c_{j}=0$, the inequality does not involve $T_{j}$ and reduces to the constraint $a_{\ell}^{\top}W_{j}\leq b_{\ell}$ on $W_{j}$ alone.
Conditioning on $W_{j}$ fixes the quantities $a_{\ell}^{\top}W_{j}$, so the constraints involving $T_{j}$ take the form
$T_{j}\leq u_{\ell}(W_{j})\quad\text{or}\quad T_{j}\geq\ell_{\ell}(W_{j}),$
for suitable functions $u_{\ell}(\cdot)$ and $\ell_{\ell}(\cdot)$ derived from the ratios above. Intersecting these finitely many half-lines yields either the empty set or a (possibly unbounded) interval. Formally, define
$a_{j}(W_{j})=\sup_{\ell:a_{\ell}^{\top}c_{j}<0}\frac{b_{\ell}-a_{\ell}^{\top}W_{j}}{a_{\ell}^{\top}c_{j}},\qquad b_{j}(W_{j})=\inf_{\ell:a_{\ell}^{\top}c_{j}>0}\frac{b_{\ell}-a_{\ell}^{\top}W_{j}}{a_{\ell}^{\top}c_{j}},$
with the convention that the supremum over an empty set is -$\infty$ and the infimum over an empty set is $\infty$. Then, on the event that all constraints with $a_{\ell}^{\top}c_{j}=0$ are satisfied by $W_{j}$, the selection event can be written as
$\mathcal{E}=\{Ay^{(\mathrm{sur})}\leq b\}=\{a_{j}(W_{j})\leq T_{j}\leq b_{j}(W_{j})\}.$
In particular, conditioning on $W_{j}$ and $\mathcal{E}$ restricts $T_{j}$ to the random interval $a_{j}(W_{j}),b_{j}(W_{j})]$.
Truncated-normal characterization.
Combining the Gaussianity of $(T_{j},W_{j})$ with the description of $\mathcal{E}$ in the $(T_{j},W_{j})$ coordinates yields the standard truncated-normal law.
Lemma B.1 (Truncated-normal law for $T_{j}$).
Work under the Gaussian surrogate model (1) and the polyhedral selection event (2). For a fixed selected feature $j\in S$, let $T_{j}=v_{j}^{\top}y^{(\mathrm{sur})}$ and $W_{j}=(I_{n}-c_{j}v_{j}^{\top})y^{(\mathrm{sur})}$ be defined as in (7), with $\theta_{j}$ and $\tau_{j}^{2}$ as above. Condition on $W_{j}$ and on the event $\mathcal{E}$ that the selection procedure outputs $S$. Then
$T_{j}\mid(W_{j},\mathcal{E})\sim\mathcal{N}(\theta_{j},\tau_{j}^{2})\text{ truncated to the interval }[a_{j}(W_{j}),b_{j}(W_{j})].$
Proof.
The joint Gaussianity of $(T_{j},W_{j})$ with $T_{j}\sim\mathcal{N}(\theta_{j},\tau_{j}^{2})$ and $T_{j}$ independent of $W_{j}$ follows from the linear transformation (7) applied to a spherical Gaussian. The preceding discussion shows that, for any fixed realization of $W_{j}$ that satisfies the constraints with $a_{\ell}^{\top}c_{j}=0$, the selection event $\mathcal{E}$ is equivalent to the requirement that $T_{j}$ lie in the interval $a_{j}(W_{j}),b_{j}(W_{j})]$. Thus the conditional law of $T_{j}$ given $(W_{j},\mathcal{E})$ is obtained by restricting the $\mathcal{N}(\theta_{j},\tau_{j}^{2})$ density to this interval and renormalizing, which is precisely a truncated-normal distribution. ∎
Lemma B.1 is the basis for the one-dimensional selective-inference map
$\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)=\bigl(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j}\bigr)$
used in the main text: under $H_{0,j}:\theta_{j}=0$ we evaluate tail probabilities in the truncated-normal law of $T_{j}$ conditional on $(W_{j},\mathcal{E})$, and we invert this family of tests to obtain selective confidence intervals for $\beta_{j}$.
B.2 Proof of selective validity
We now justify the selective validity statement in Proposition 4.2 using the truncated-normal characterization from Lemma B.1. Throughout this subsection we work under the Gaussian surrogate model (1) and the polyhedral selection event (2), and we condition on $(f,X,S_{0})$ as in Lemma 4.1. For a fixed selected feature $j\in S$, recall that
$T_{j}=v_{j}^{\top}y^{(\mathrm{sur})},\qquad\theta_{j}=v_{j}^{\top}\mu,\qquad\tau_{j}^{2}=\sigma^{2}\|v_{j}\|_{2}^{2},$
and that Lemma B.1 implies
$T_{j}\mid(W_{j},\mathcal{E})\sim\mathcal{N}(\theta_{j},\tau_{j}^{2})\text{ truncated to }[a_{j}(W_{j}),b_{j}(W_{j})].$
Conditional tests based on truncated-normal quantiles.
Fix a realization $W_{j}=w$ for which the selection event is feasible, so that Lemma B.1 yields a truncated-normal law for $T_{j}\mid(W_{j}=w,\mathcal{E})$. For notational ease, write
$a(w),b(w)]=[a_{j}(w),b_{j}(w)]$
for the corresponding truncation interval, and denote by
$F_{0,w}(t)=\mathbb{P}\bigl(T_{j}\leq t\,\big|\,W_{j}=w,\mathcal{E},H_{0,j}:\theta_{j}=0\bigr)$
the cumulative distribution function of $T_{j}$ under the null hypothesis $\theta_{j}=0$, conditional on $(W_{j}=w,\mathcal{E})$. By Lemma B.1, $F_{0,w}$ is the CDF of a univariate Gaussian with mean zero and variance $\tau_{j}^{2}$, truncated to $[a(w),b(w)]$.
For concreteness, consider a one-sided test of $H_{0,j}:\theta_{j}=0$ against the alternative $\theta_{j}>0$. For each $\alpha\in(0,1)$ and each fixed $w$, define the critical value $c_{\alpha}(w)$ as the upper $(1-\alpha)$ quantile of this truncated-normal law under the null:
$c_{\alpha}(w)=F_{0,w}^{-1}(1-\alpha),$
so that
$\mathbb{P}\bigl(T_{j}>c_{\alpha}(w)\,\big|\,W_{j}=w,\mathcal{E},H_{0,j}\bigr)=\alpha.$
The corresponding critical region is
$C_{\alpha}(w)=\{t\in[a(w),b(w)]:t>c_{\alpha}(w)\}.$
By construction, for every fixed $w$ and every $\alpha\in(0,1)$,
$\mathbb{P}\bigl(T_{j}\in C_{\alpha}(w)\,\big|\,W_{j}=w,\mathcal{E},H_{0,j}\bigr)=\alpha.$ (8)
We define the selective $p$-value as the smallest significance level at which the observed $T_{j}$ would fall into the corresponding critical region:
$p_{j}^{\mathrm{sel}}=\inf\bigl\{\alpha\in(0,1):T_{j}\in C_{\alpha}(W_{j})\bigr\}.$ (9)
This is the usual definition of a $p$-value via a nested family of tests; for a two-sided alternative one can replace $C_{\alpha}(w)$ by a symmetric critical region in both tails, but the argument below is unchanged.
Conditional type-I error control.
We now verify that $p_{j}^{\mathrm{sel}}$ is conditionally valid given $(f,X,S_{0},j\in S)$, as claimed in Proposition 4.2. It suffices to establish the bound conditional on $W_{j}$ and $\mathcal{E}$, and then integrate over $W_{j}$.
Fix any $\alpha\in(0,1)$. By the definition (9), the event $\{p_{j}^{\mathrm{sel}}\leq\alpha\}$ is exactly the event that $T_{j}$ falls into the $\alpha$-level critical region:
$\{p_{j}^{\mathrm{sel}}\leq\alpha\}=\{T_{j}\in C_{\alpha}(W_{j})\}.$
Therefore, under $H_{0,j}$ and conditional on $(W_{j}=w,\mathcal{E})$, (8) implies
$\mathbb{P}\bigl(p_{j}^{\mathrm{sel}}\leq\alpha\,\big|\,W_{j}=w,\mathcal{E},H_{0,j}\bigr)=\mathbb{P}\bigl(T_{j}\in C_{\alpha}(w)\,\big|\,W_{j}=w,\mathcal{E},H_{0,j}\bigr)=\alpha.$
Taking expectations with respect to the conditional law of $W_{j}$ given $(\mathcal{E},f,X,S_{0})$, we obtain
$\mathbb{P}\bigl(p_{j}^{\mathrm{sel}}\leq\alpha\,\big|\,\mathcal{E},f,X,S_{0},H_{0,j}\bigr)=\mathbb{E}\bigl[\mathbb{P}(p_{j}^{\mathrm{sel}}\leq\alpha\mid W_{j},\mathcal{E},f,X,S_{0},H_{0,j})\bigr]=\alpha.$
In particular,
$\mathbb{P}\bigl(p_{j}^{\mathrm{sel}}\leq\alpha\,\big|\,f,X,S_{0},j\in S,H_{0,j}\bigr)\leq\alpha,$
which is the desired conditional type-I error control.
Selective confidence intervals by test inversion.
We now show that the confidence intervals obtained by inverting the truncated-normal tests achieve the advertised conditional coverage. For each candidate value $\theta\in\mathbb{R}$, define the null model $H_{j}(\theta):\theta_{j}=\theta$ and let $F_{\theta,w}$ denote the CDF of $T_{j}$ under $H_{j}(\theta)$, conditional on $(W_{j}=w,\mathcal{E})$. For each $\theta$ and each $\alpha\in(0,1)$, we can construct an $\alpha$-level test of $H_{j}(\theta)$ based on the truncated-normal quantiles under $F_{\theta,w}$; denote the corresponding critical region by $C_{\alpha}(w;\theta)$. By construction,
$\mathbb{P}\bigl(T_{j}\in C_{\alpha}(W_{j};\theta)\,\big|\,W_{j},\mathcal{E},H_{j}(\theta)\bigr)=\alpha.$
The $(1-\alpha)$ selective confidence interval for $\theta_{j}$ is then defined by test inversion:
$\mathrm{CI}_{j}=\Bigl\{\theta\in\mathbb{R}:T_{j}\notin C_{\alpha}(W_{j};\theta)\Bigr\}.$ (10)
Equivalently, $\mathrm{CI}_{j}$ is the set of parameter values that would not be rejected at level $\alpha$ given the observed $(T_{j},W_{j})$. Let $\theta_{j}^{\star}$ denote the true value of $\theta_{j}$. Using (10) with $\theta=\theta_{j}^{\star}$ and the fact that the corresponding test has level $\alpha$, we obtain
$\mathbb{P}\bigl(\theta_{j}^{\star}\in\mathrm{CI}_{j}\,\big|\,W_{j},\mathcal{E},H_{j}(\theta_{j}^{\star})\bigr)=\mathbb{P}\bigl(T_{j}\notin C_{\alpha}(W_{j};\theta_{j}^{\star})\,\big|\,W_{j},\mathcal{E},H_{j}(\theta_{j}^{\star})\bigr)=1-\alpha.$
Integrating over the conditional law of $W_{j}$ given $(\mathcal{E},f,X,S_{0})$ yields
$\mathbb{P}\bigl(\theta_{j}^{\star}\in\mathrm{CI}_{j}\,\big|\,\mathcal{E},f,X,S_{0}\bigr)=1-\alpha,$
and hence
$\mathbb{P}\bigl(\beta_{j}\in\mathrm{CI}_{j}\,\big|\,f,X,S_{0},j\in S\bigr)\geq 1-\alpha,$
where we use that $\theta_{j}=v_{j}^{\top}X_{S_{0}}\beta$ is the linear functional of $\beta$ corresponding to feature $j$ in the surrogate model. This establishes the conditional coverage claim in Proposition 4.2.
Appendix C Implementation and Computational Details
We summarize the main implementation details for the experiments in Section 5; full code is provided as supplementary material.
Datasets and splits.
We use four standard tabular regression benchmarks: Concrete (compressive strength prediction with $p=8$ physical inputs), AirQuality (urban air-pollution measurements with multiple gas sensors), EnergyEfficiency (building energy loads from eight geometric and material attributes), and kin8nm (kinematic response as a function of eight physical variables). For Concrete, we load the Concrete Compressive Strength data from OpenML (with the UCI file as a fallback), treat compressive strength as the target, and use the remaining eight variables as continuous features with shortened column names. For AirQuality, we load the AirQualityUCI data, drop date/time columns and rows containing the sentinel code 200-200, and use CO(GT) as the target with all remaining numerical variables as features. Unless otherwise stated, for each repetition we draw a random split with 80% training data and 20% test data. The test split is held out from all feature-selection and inference procedures and is used only to evaluate predictive fidelity. Within the 80% training portion, the $\phi$-test further splits the data into a selection half and an inference half of equal size.
Backbones and surrogate targets.
On each dataset we fit two black-box predictors on the training split: a gradient-boosted tree model (XGBRegressor with a standard configuration) and a feed-forward multilayer perceptron (two hidden layers with ReLU activations, trained with Adam and early stopping). For the MLP we standardize all input features via a StandardScaler–MLP pipeline. Given a trained backbone $f$, we define the surrogate target $y^{(\mathrm{sur})}=f(X)$ on the training covariates and evaluate the backbone performance on the held-out test set via the coefficient of determination $R^{2}_{\mathrm{full}}$.
SHAP computation and global scores.
For tree backbones we compute TreeSHAP values using shap.TreeExplainer; for MLP backbones we use KernelSHAP with a small background sample and a capped number of evaluation points. In both cases we compute SHAP values on the selection half of the training data and define global scores $I_{j}=\mathbb{E}|\phi_{j}(x)|$ as the mean absolute SHAP value over this subset. The held-out test split is never used for SHAP computation or feature selection. On the three eight-feature datasets (Concrete, EnergyEfficiency, and kin8nm) we screen to the top $M=7$ features by $I_{j}$, and on AirQuality to the top $M=10$; the remaining features are discarded for the $\phi$-test but remain available to the baselines.
Implementation of the $\phi$-test.
Stage 1 (Subsection 3.1) forms the screened design $X_{S_{0}}$ by retaining the top-$M$ features by global SHAP score. Stage 2 (Subsection 3.2) applies LARS-lasso (lars_path with the lasso option) to the standardized selection-half design on $S_{0}$ and selects the first $K=5$ variables that enter the path. Stage 3 (Subsection 3.3) fits ordinary least squares for $y^{(\mathrm{sur})}$ on the selected features $S$. We compute two sets of regression summaries: (i) a split-sample fit using only the inference half (independent of the selection step), which yields coefficients, standard errors, $t$-statistics, $p$-values, and confidence intervals that are classically valid and can be viewed as a conservative special case of the selective-inference framework in Section 4; and (ii) a naive fit using the full training data, which illustrates the effect of ignoring selection. The global feature-importance table for the $\phi$-test combines the SHAP scores, the selective summaries, and the naive summaries into a single per-feature table and adds a residual row that aggregates the SHAP mass of unselected features.
Baselines and evaluation metrics.
All four Shapley-based baselines share the same backbones, surrogate targets, and SHAP values as the $\phi$-test. SHAP-TopK selects the top $K$ features by global score $I_{j}$. SPVIM-Boot bootstraps rows of the SHAP matrix to obtain a standard error for $I_{j}$ and selects features whose normal-approximation $p$-values are below $0.05$. SHAP-HT performs a one-sample $t$-test on $\phi_{j}(x)$ for each feature with Bonferroni correction. StableSHAP repeatedly bootstraps the SHAP matrix, records the top-$K$ set in each resample, and selects features that appear with frequency at least $0.7$. Given a selected set $S$, we fit an OLS surrogate on $X_{S}$ and report the test performance ratio $100\times R^{2}_{\mathrm{sel}}/R^{2}_{\mathrm{full}}$, the number of selected features, and a Jaccard stability index computed over repeated train–test splits. Backbone robustness is measured by the Jaccard index between the selected sets obtained from the XGBoost and MLP backbones on a common split. A small ablation on Concrete compares four variants of the Stage 2 (Subsection 3.2) selection step (with and without SHAP screening, and with LARS versus forward stepwise selection) using the same fidelity, sparsity, stability, and backbone-robustness metrics.